# MMedAgent-RL: Optimizing Multi-Agent Collaboration for Multimodal Medical Reasoning

### v2

@@ E.2 Scaling Up the Model Parameters to 7B
GPT-4o $\rightarrow$ Qwen2.5-VL-3B 61.8 60.3 49.3 57.1 52.0 46.2 49.6
GPT-4o $\rightarrow$ Qwen2.5-VL-3B+SFT 61.4 61.1 54.3 59.0 54.1 44.8 49.5
MMedAgent-RL (3B) 67.9 69.8 69.2 69.0 61.5 57.5 59.5
MMedAgent-RL (7B) 70.2 75.3 72.5 72.7 68.9 66.4 67.7
E.3 Detailed Results
Traditional Medical Imaging Evaluation.
Table 7 presents the accuracy of various models across five major medical imaging modalities in the OmniMedVQA benchmark. Our model (MMedAgent-RL) demonstrates strong generalization across all categories, achieving an average accuracy of 61.5%, significantly outperforming previous state-of-the-art models including LLaVA-v1.6-34B (58.7%) and Qwen2.5-VL-7B (60.8%).
@@ E.3 Detailed Results
Specifically, our method achieves 66.2% on microscopy images, indicating robust capability in processing fine-grained, high-resolution visual data typical of pathology slides. On MRI and CT modalities, MMedAgent-RLreaches 63.4% and 61.3%, respectively, outperforming strong baselines such as LLaVA-v1.6-34B and Yi-VL-34B by a wide margin. These results show that our model captures both structural and soft-tissue anatomical details effectively. In X-Ray, our method maintains competitive performance (65.0%) compared to high-performing models like HuatuoGPT-Vision-7B (80.3%), while achieving the highest accuracy on Ultrasound (51.7%) among all models, demonstrating robustness in handling noisy, low-contrast imaging modalities.
MMMU Health & Medicine Track.
In Table 8, our model again establishes new performance standards, achieving 57.5% overall accuracy on the MMMU Health & Medicine test set. Compared to existing large models such as Qwen2.5-VL-7B (56.6%) and HuatuoGPT-Vision-7B (51.0%), MMedAgent-RLdemonstrates clear advantages.
@@
Notably, our model excels across all five sub-domains: scoring 64.7% in Basic Medical Science (BMS), 63.0% in Clinical Medicine (CM), 53.0% in Diagnostics and Laboratory Medicine (DLM), 56.3% in Pharmacy (P), and 50.5% in Public Health (PH). These results reflect a well-rounded capability across both foundational scientific understanding and applied clinical knowledge. In particular, performance in CM and P shows substantial improvement over single-agent baselines, suggesting that our model benefits from enhanced reasoning and domain transfer.
Taken together, these results confirm the effectiveness of our approach in both imaging-based and knowledge-based medical VQA settings, and highlight the potential of our method as a comprehensive solution for multimodal medical understanding.
@@
GPT-4o $\rightarrow$ Qwen2.5-VL-3B 57.3 50.2 60.6 65.0 26.7 52.0
GPT-4o $\rightarrow$ Qwen2.5-VL-3B+SFT 56.7 51.1 58.9 70.1 27.5 52.9
MMedAgent-RL (3B) 61.3 63.4 66.2 65.0 51.7 61.5
@@
GPT-4o $\rightarrow$ Qwen2.5-VL-3B 50.0 53.3 36.7 55.6 36.7 46.2
GPT-4o $\rightarrow$ Qwen2.5-VL-3B+SFT 50.0 50.0 26.7 55.6 43.3 44.8
MMedAgent-RL (3B) 64.7 63.0 53.0 56.3 50.5 57.5
E.4 More Cases
To further demonstrate the robustness and versatility of our proposed model in multimodal medical applications, we present additional representative cases in Figure 8, Figure 9, Figure 10, Figure 11, Figure 12 and Figure 13. These examples encompass various clinical scenarios and imaging modalities, providing a comprehensive illustration of the model’s ability to effectively integrate and interpret diverse types of medical data.
Figure 8: Case 1 (OmniMedVQA) & Case2 (PathVQA). In the model responses, blue text represents the process of reasoning about relevant medical knowledge based on the question, and red text represents the analysis of the answer provided by the specialists.
Figure 9: Case 3 (OmniMedVQA). In the model responses, blue text represents the process of reasoning about relevant medical knowledge based on the question, and red text represents the analysis of the answer provided by the specialists.
Figure 10: Case 4 (PathVQA). In the model responses, blue text represents the process of reasoning about relevant medical knowledge based on the question, and red text represents the analysis of the answer provided by the specialists.
Figure 11: Case 5 (VQA-RAD) & Case 6 (VQA-RAD). In the model responses, blue text represents the process of reasoning about relevant medical knowledge based on the question, and red text represents the analysis of the answer provided by the specialists.
Figure 12: Case 7 (PathVQA) & Case 8 (MMMU). In the model responses, blue text represents the process of reasoning about relevant medical knowledge based on the question, and red text represents the analysis of the answer provided by the specialists.
Figure 13: Case 9 (MMMU). In the model responses, blue text represents the process of reasoning about relevant medical knowledge based on the question, and red text represents the analysis of the answer provided by the specialists.

### v3

@@ G.1 Comparison with Supervised Fine-Tuning (SFT)
AFlow 67.3 68.9 66.4 67.5 59.6 53.6 56.6
GPT-4o $\rightarrow$ Qwen2.5-VL-7B 62.5 63.9 53.2 59.9 56.4 50.7 53.6
GPT-4o $\rightarrow$ Qwen2.5-VL-7B+SFT w/o reasoning 65.5 66.5 61.4 64.5 60.9 57.8 62.4
GPT-4o $\rightarrow$ Qwen2.5-VL-7B+SFT w/ reasoning 68.8 68.4 63.7 67.0 62.0 59.4 64.5
MMedAgent-RL (7B) 71.5+10% 76.2+12% 72.3+12% 73.3+11% 73.3+13% 71.9+15% 72.6+14%
G.2 Different Specialists
The effectiveness of different specialist compositions within our framework is detailed in Table 6. The results unequivocally show that multi-agent collaboration substantially outperforms single-agent baselines. Our premier configuration, 3*OpenAI-o3 (OpenAI, 2025), achieved a top average score of 73.0, far exceeding the best-performing baseline, GPT-4o (68.8). Crucially, this performance gain stems from the synergistic integration of multiple experts. The analysis also underscores the importance of specialist diversity. Heterogeneous teams combining different models (e.g., 2*OpenAI-o3+1*HuatuoGPT at 71.3) proved highly effective, demonstrating that fusing complementary knowledge enhances diagnostic robustness. This confirms that our framework’s core strength lies in its ability to dynamically orchestrate collaboration among a diverse team of high-quality specialists to achieve superior decision-making.
Table 6: The comparison with different specialists.
Model VQA-RAD SLAKE PathVQA OmniMedVQA MMMU Avg.
GPT-4o 61.0 75.5 69.4 68.5 69.7 68.8
Qwen2.5-VL-7B 61.8 64.7 60.5 60.8 56.6 60.9
HuatuoGPT-7B 63.0 77.2 58.7 74.6 51.0 64.9
MMedAgent-RL
3*OpenAI-o3 71.5 76.2 72.3 73.3 71.9 73.04
3*GPT-4o 70.4 75.2 72.7 69.1 67.1 70.9
1*GPT-4o 63.2 65.9 62.8 63.4 59.0 62.86
1*HuatuoGPT-Vision-7B 63.4 67.2 60.8 64.9 54.1 62.1
1*Qwen2.5-VL-7B 62.4 63.2 61.5 61.5 55.1 60.7
2*GPT-4o+1*Qwen2.5-VL-7B 70.0 74.6 71.3 66.0 64.0 69.2
2*GPT-4o+1*HuatuoGPT-Vision-7B 68.5 75.0 71.7 68.0 62.0 69.0
2*OpenAI-o3+1*HuatuoGPT-Vision-7B 69.9 75.8 73.0 70.0 68.0 71.3
2*OpenAI-o3+1*Qwen2.5-VL-7B 71.1 75.4 71.8 69.1 69.2 71.3
1*Qwen2.5-VL-7B+1*GPT-4o 68.9 73.8 70.6 66.1 63.2 68.5
1*HuatuoGPT-Vision-7B+1*OpenAI-o3 69.0 74.7 72.3 68.8 68.3 70.6
3*HuatuoGPT-Vision-7B 65.8 78.2 61.3 73.8 50.1 65.8
G.3 Ablation Analysis
G.3.1 Performance of Triage Doctor
Table 7: The performance of triage doctor.
Model VQA-RAD SLAKE PathVQA
Qwen2.5-VL-3B 95.62 92.16 77.53
Qwen2.5-VL-7B 96.21 94.41 80.58
MMedAgent-RL 99.98 99.94 99.06
G.3.2 KL Divergence Coefficient
Figure 6: Ablation of KL divergence coefficient.
G.4 Comparison with Test-Time Scaling Methods
We conducted additional experiments using test-time scaling techniques, specifically Majority Voting and Self-Consistency (Wang et al., ), on both Qwen2.5-VL-7B and GPT-4o. For the experimental setup, Majority Voting involved sampling $N=3$ outputs and selecting the most frequent answer as the final prediction. Self-Consistency sampled a set of diverse reasoning paths rather than relying on the greedy path, subsequently selecting the most consistent answer by marginalizing over the sampled paths. As shown in Table 8, test-time scaling improved the performance of Qwen2.5-VL-7B by 2.6% and GPT-4o by 1.6% across the five datasets. However, despite these improvements, a performance gap remains compared to our multi-agent framework. Our method outperforms these enhanced baselines by 18%, demonstrating the necessity and effectiveness of our proposed triage and multi-expert pipeline.
Table 8: Performance comparison with test-time scaling baselines.
Method VQA-RAD SLAKE PathVQA OmniMedVQA MMMU-Med Overall
Qwen2.5-VL-7B 61.8 64.7 60.5 60.8 56.6 60.9
Qwen2.5-VL-7B + Majority Voting 63.7 65.4 61.5 62.5 57.0 62.0
Qwen2.5-VL-7B + Self-consistency 63.8 65.3 62.4 64.5 59.1 63.0
GPT-4o 61.0 75.5 69.4 68.5 69.7 68.8
GPT-4o + Majority Voting 62.2 75.9 70.3 69.4 70.3 69.6
GPT-4o + Self-consistency 63.6 75.4 70.8 69.1 71.5 70.1
MMedAgent-RL 71.5 76.2 72.3 73.3 71.9 73.0
G.5 Quantify the Performance Gain from Routing and Aggregation
To strictly quantify the model’s ability to correct mistakes rather than simply route to a specialist, we conducted a targeted evaluation on hard samples where all three specialists provided incorrect answers. We selected 200 such samples from PathVQA and 200 from OmniMedVQA to evaluate the performance of the Base Model (with and without expert knowledge), the SFT baseline, and our MMedAgent-RL. Since every specialist is incorrect, any routing mechanism would yield 0% accuracy. Therefore, the performance on this subset is strictly attributable to the model’s ability to perform intrinsic correction. As shown in Table 9, our method achieves substantial improvements. For instance, on the OOD OmniMedVQA hard subset, MMedAgent-RL improves accuracy to 23.0%. We attribute this capability to the following logic: The extremely low accuracy of the Base Model w/o expert knowledge (4.5% on PathVQA and 2.0% on OmniMedVQA) confirms that the model lacks the intrinsic parametric knowledge to solve these hard cases independently. Although the specialists’ final answers were wrong, their reasoning processes likely contained partial truths or valid medical context. While the SFT baseline struggles to utilize this conflicting information (often hallucinating along with the experts), MMedAgent-RL has learned via RL to critically synthesize these valid reasoning fragments, correcting the final conclusion rather than simply aggregating the errors.
Table 9: Correctness ratio (accuracy on hard samples where all specialists failed).
Setting / Model PathVQA OmniMedVQA
w/o expert knowledge
   Base Model 4.5% 2.0%
w/ expert knowledge
   Base Model 5.5% 3.5%
   SFT 10.0% 7.5%
   MMedAgent-RL 26.5% 23.0%
G.6 Detailed Ablation on Triage Agent
In Table 2, “w/o Triage” refers to using the original base model to perform the routing (triage) task directly, without specific fine-tuning for this role. Following your suggestions, we have expanded our comparison in Table 10. We introduced two new settings to test the necessity of the framework: 1) Random Routing: Assigning queries to specialists randomly to isolate the benefit of the specialists themselves. 2) Single Model w/o Routing: Using the Qwen2.5-VL-7B base model directly, and enhancing it with Majority Voting (diverse sampling) as requested.
As shown in Table 10, incorporating a routing mechanism leads to significant performance gains. Notably, even Random Routing generally outperforms the single model equipped with Majority Voting. Furthermore, our proposed MMedAgent-RL significantly outperforms all baselines, confirming that a dedicated triage-and-expert pipeline provides advantages that cannot be achieved by simple diverse sampling or random assignment.
Comparison with SFT. The original “w/o Triage” ablation was insufficient to justify the specific choice of RL over simpler methods. To address this, we conducted a detailed comparison between our GRPO-optimized Triage Agent and standard Supervised Fine-Tuning (SFT) baselines. As shown in Table 11, we evaluated two SFT configurations: 1) SFT (Standard): Fine-tuned on direct Question-Department pairs. 2) SFT (w/ Reasoning): Fine-tuned using reasoning traces distilled from Qwen2.5-VL-32B to simulate a more capable classifier. While SFT significantly improves performance over the base model, GRPO still outperforms the best SFT baseline across all datasets. Crucially, this advantage is most pronounced on Out-of-Distribution (OOD) datasets. For instance, on MMMU-Med, GRPO outperforms “SFT w/ Reasoning” by +4.1% (71.9 vs. 67.8), and on OmniMedVQA by +2.5% (73.3 vs. 70.8). This indicates that while SFT can achieve high accuracy on standard distributions, the reinforcement learning process (GRPO) enables the Triage Agent to generalize better to complex, unseen scenarios by learning from the reward signal of the downstream reasoning success, rather than just mimicking a static label.
Table 10: Performance comparison with or without routing.
Method VQA-RAD SLAKE PathVQA OmniMedVQA MMMU-Med Overall
w/ triage
   MMedAgent-RL (zero-shot) 66.3 69.9 67.2 66.2 59.3 65.8
   MMedAgent (random triage) 63.4 66.7 64.7 67.8 58.2 64.2
   MMedAgent-RL (fine-tuned triage) 71.5 76.2 72.3 73.3 71.9 73.0
w/o triage
   Qwen2.5-VL-7B 61.8 64.7 60.5 60.8 56.6 60.9
   + Majority Voting 63.7 65.4 61.5 62.5 57.0 62.0
Table 11: Performance comparison of Triage Agent training strategies.
Method VQA-RAD SLAKE PathVQA OmniMedVQA MMMU-Med
Base model 66.3 69.9 67.2 66.2 59.3
+ SFT w/o reasoning process 69.4 75.0 70.5 70.1 66.7
+ SFT w/ reasoning process 70.2 75.9 71.0 70.8 67.8
+ GRPO 71.5 76.2 72.3 73.3 71.9
G.7 Detailed Ablation on Progressively Adding Components
To quantify the contribution of each component, we have conducted a progressive evaluation as shown in Table 12. Regarding the order of ablation, we formulated the progression as Base → Multi-expert → Triage. From an architectural perspective, the Triage module depends on the existence of a candidate pool of experts to perform routing. Thus, we first introduce the Specialists (Multi-expert) to build the capability pool, and subsequently add the Triage module to manage and utilize these experts efficiently. Starting with the Qwen2.5-VL-7B baseline, we observed the following trends:
  • + Majority Voting: Provides a marginal improvement, indicating that simple test-time scaling has limits.
  • + Specialists: Integrating domain-specific experts (with a base model as the router) yields further gains, surpassing the single model with voting.
  • + Triage: Introducing the learned Triage module significantly improves the effective utilization of specialists.
  • + Curriculum RL: Finally, applying our Curriculum RL strategy provides the most substantial performance leap, demonstrating that optimizing the collaboration between the triage and specialist agents is critical for complex medical reasoning.
Table 12: Performance progressively adding components.
Method VQA-RAD SLAKE PathVQA OmniMedVQA MMMU-Med
Qwen2.5-VL-7B (Base) 61.8 64.7 60.5 60.8 56.6
+ Specialists (Base Model) 64.5 66.9 63.2 63.4 60.7
+ Triage 65.7 68.4 64.4 64.8 62.6
+ Curriculum RL 71.5 76.2 72.3 73.3 71.9
G.8 Comparison on Triage Agent with Different Settings
As shown in Table 13, we quantitatively evaluated the impact of different training stages on the triage agent’s performance. We compared the Base Model, SFT (trained on direct question-answer pairs without reasoning), SFT with Reasoning (trained on reasoning traces distilled from Qwen2.5-VL-32B), and our final GRPO-optimized model. The results demonstrate that the triage agent trained with GRPO yields the highest performance. Notably, this improvement is most significant on the two Out-of-Distribution (OOD) datasets, i.e., OmniMedVQA and MMMU-Med, confirming that the reasoning capabilities reinforced by GRPO are crucial for generalization.
Table 13: Performance with triage agent with different settings.
Method VQA-RAD SLAKE PathVQA OmniMedVQA MMMU-Med Overall
Base model 66.3 69.9 67.2 66.2 59.3 65.8
+ SFT w/o reasoning process 69.4 75.0 70.5 70.1 66.7 70.3
+ SFT w/ reasoning process 70.2 75.9 71.0 70.8 67.8 71.1
+ GRPO 71.5 76.2 72.3 73.3 71.9 73.0
G.9 Comparison on GP Update Strategies
In the paper, the two GP agents (the Triage Agent and the Attending Physician) are updated independently. This design choice was primarily made to ensure training stability and to decouple the training process from potential API failures or latency when querying the external OpenAI-based specialists. Specifically, our procedure is as follows: 1) Triage Optimization: We first optimize the Triage Agent using image-modality QA pairs to ensure accurate department routing. 2) Data Preparation: We classify the training data based on these departments and invoke the OpenAI API (acting as specialists) to generate expert knowledge offline. 3) Attending GP Training: Finally, we use these pre-generated expert trajectories to train the Attending GP (Qwen2.5-VL) via Reinforcement Learning.
We also implemented an end-to-end online framework where both GPs are updated simultaneously. As shown in Table 14, the performance difference between the two settings is negligible. This confirms that our decoupled training strategy is valid and yields results consistent with a fully end-to-end approach while remaining more computationally efficient and stable.
Table 14: Performance comparison of GP update strategies.
Strategy VQA-RAD SLAKE PathVQA OmniMedVQA MMMU-Med Overall
Independent 71.5 76.2 72.3 73.3 71.9 73.0
Simultaneous 71.3 76.5 72.4 73.6 71.6 73.1
G.10 Framework Transferability
To demonstrate the transferability of our framework, we conducted additional experiments using InternVL2.5-Instruct-8B (Chen et al., 2024b) as an alternative base model. As shown in Table 15, our method yields consistent and significant improvements across all datasets, regardless of the backbone architecture. On OmniMedVQA, the InternVL-based agent achieved a remarkable score of 82.4%, surpassing the performance of the Qwen-based version. Even on datasets where the base InternVL model struggled (e.g., PathVQA, where the base score was only 42.3%), our framework provided a massive performance boost of +26.1% (reaching 68.4%). These results confirm that our pipeline is model-agnostic and can effectively enhance the reasoning capabilities of diverse multimodal LLMs.
Table 15: Assessment of framework transferability.
Model VQA-RAD SLAKE PathVQA OmniMedVQA MMMU-Med
Qwen2.5-VL-7B 61.8 64.7 60.5 60.8 56.6
MMedAgent-RL (Qwen2.5-VL-7B) 71.5 76.2 72.3 73.3 71.9
InternVL2.5-8B 58.6 68.6 42.3 76.5 51.4
MMedAgent-RL (InternVL2.5-8B) 70.2 78.9 68.4 82.4 64.7
G.11 Baselines with Trained Aggregator
We implemented two new baselines where GPT-4o first samples $N=3$ diverse outputs per query ($T=1.0$), and a Qwen2.5-VL-7B model is then trained on the training set (same as data we used) to act as an aggregator that selects or synthesizes the final answer from these candidates. We developed both an SFT Aggregator (via Supervised Fine-Tuning) and a GRPO Aggregator (via Group Relative Policy Optimization) to ensure a robust comparison. As shown in Table 16, MMedAgent-RL still maintains a significant performance lead. We observed that directly training an aggregator yields limited gains, particularly when facing inconsistent candidate answers, i.e., a challenge that directly motivated our proposed curriculum learning-guided RL strategy. This confirms that our method’s effectiveness stems from the process-level collaboration of specialized agents, which captures domain-specific nuances that cannot be replicated by simply aggregating generalist outputs.
Table 16: Performance comparison with GPT-4o+an aggregator (based on Qwen2.5-VL 7B) baselines.
Model VQA-RAD SLAKE PathVQA OmniMedVQA MMMU-Med Overall
Qwen2.5-VL-7B (Base) 61.8 64.7 60.5 60.8 56.6 60.9
GPT-4o 61.0 75.5 69.4 68.5 69.7 68.8
GPT-4o+Qwen2.5-VL-7B (SFT Fine-tuned) 65.8 69.4 62.9 60.3 65.3 64.7
GPT-4o+Qwen2.5-VL-7B (GRPO Fine-tuned) 67.3 70.8 63.8 61.8 64.6 65.7
MMedAgent-RL 71.5 76.2 72.3 73.3 71.9 73.0
G.12 Detailed Results
@@
Traditional Medical Imaging Evaluation. Table 17 presents the accuracy of various models across five major medical imaging modalities in the OmniMedVQA benchmark. Our model (MMedAgent-RL) demonstrates strong generalization across all categories, achieving an average accuracy of 73.3%, significantly outperforming previous state-of-the-art models including LLaVA-v1.6-34B (58.7%) and Qwen2.5-VL-7B (60.8%). Specifically, our method achieves 76% on microscopy images, indicating robust capability in processing fine-grained, high-resolution visual data typical of pathology slides. On MRI and CT modalities, MMedAgent-RL reaches 72% and 65%, respectively, outperforming strong baselines such as LLaVA-v1.6-34B and Yi-VL-34B by a wide margin. These results show that our model captures both structural and soft-tissue anatomical details effectively. In X-Ray, our method maintains competitive performance (78.8%) compared to high-performing models like HuatuoGPT-Vision-7B (80.3%), while achieving the highest accuracy on Ultrasound (75%) among all models, demonstrating robustness in handling noisy, low-contrast imaging modalities.
@@
MMMU Health & Medicine Track. In Table 18, our model again establishes new performance standards, achieving 71.9% overall accuracy on the MMMU Health & Medicine test set. Compared to existing large models such as Qwen2.5-VL-7B (56.6%) and HuatuoGPT-Vision-7B (51.0%), MMedAgent-RLdemonstrates clear advantages. Notably, our model excels across all five sub-domains: scoring 75% in Basic Medical Science (BMS), 78% in Clinical Medicine (CM), 65% in Diagnostics and Laboratory Medicine (DLM), 70% in Pharmacy (P), and 71.5% in Public Health (PH). These results reflect a well-rounded capability across both foundational scientific understanding and applied clinical knowledge. In particular, performance in CM and P shows substantial improvement over single-agent baselines, suggesting that our model benefits from enhanced reasoning and domain transfer. Taken together, these results confirm the effectiveness of our approach in both imaging-based and knowledge-based medical VQA settings, and highlight the potential of our method as a comprehensive solution for multimodal medical understanding.
@@
AFlow 59.5 62.0 63.2 61.7 51.6 59.6
MMedAgent-RL (7B) 64.6 71.7 76.0 78.8 75.4 73.3
@@
MMedAgent-RL (7B) 75.0 78.0 65.0 70.0 71.5 71.9
G.13 More Cases
To further demonstrate the robustness and versatility of our proposed model in multimodal medical applications, we present additional representative cases in Figure 7, Figure 8, Figure 9 and Figure 10. These examples encompass various clinical scenarios and imaging modalities, providing a comprehensive illustration of the model’s ability to effectively integrate and interpret diverse types of medical data.
Figure 7: Case 1 (CT scan of foot): The model identifies plantar fascia pathology from a CT image, requiring anatomical knowledge of soft tissue structures in the foot and recognition of characteristic changes consistent with plantar fasciitis. Case 2 (Protein structure diagram): A structural biology reasoning task based on a protein diagram, where the absence of distinct subunit boundaries leads to the conclusion that the protein does not exhibit quaternary structure—showcasing visual-structural reasoning in molecular biology.
Figure 8: Case 3 (Ultrasound image): The model interprets an ultrasound to identify the pancreas, integrating grayscale image features with organ shape and texture—an example of soft tissue recognition using sonography. Case 4 (Brain MRI with question about vestibular system): This task highlights the model’s ability to identify mismatches between the image modality (brain MRI) and the clinical question (vestibular system), and to reason that facial sensation loss is not explainable from the given scan.
Figure 9: Case 5 (Abdominal CT image – polysplenia diagnosis): The task involves recognizing abnormal organ positioning indicative of polysplenia. This case highlights anatomical reasoning through CT imaging, requiring recognition of fused or duplicated spleens and an understanding of associated syndromic features. Case 6 (Histopathology slide – adenoma diagnosis): A histological image is used to assess the presence of malignant features. The model correctly distinguishes normal plantar fascia morphology from pathological adenoma patterns, demonstrating reasoning in pathology image interpretation.
Figure 10: Case 7 (Pie chart comparison – elemental composition): This case blends image data (a pie chart of Earth’s element composition) with textual biochemical data (composition of the human body). The reasoning required crosses domains, comparing datasets to identify overlapping elements, exemplifying multimodal cross-referencing and synthesis.